If a man runs for an hour and a half at 8 miles an hour due south; and then he continues south on a bus traveling at 55 miles an hour, for 24 minutes. How far did he travel?
Answer #1
Running: (1 1/2) * 8 = ?

Bus: 24/60 = 2/5 * 55 = ?
answered by: slimstud
Answer #2
Running
1hour = 8 miles
1.5 hors = 8 x1.5 = 12 miles

Bus
60 minutes = 55 miles
24 mins = ?
24/60 x 55 = 22miles

total = 12 + 22 =34 miles.
answered by: Aim